# SoK: Cybersecurity Assessment of Humanoid Ecosystem

### v1

@@ 4.2 A Layer-by-Layer Analysis of Threats and Defenses
Physical Layer — hardware, power, and safety baseline. The physical layer represents the most critical point of potential system compromise. Runtime attacks like firmware reflashing (P-A1) can embed persistent malware via debug interfaces, bypassing all software controls. Sensor bias (P-A2) distorts physical inputs, inducing silent drift in downstream systems. Since these attacks occur below the software stack, conventional detection mechanisms often miss them. While mitigations like secure boot (P-D1) and tamper evidence (P-D4) provide strong safeguards, their inconsistent adoption leaves foundational vulnerabilities exposed.
@@
Sensing and Perception Layer — Integrity Control Point for Environmental Input. The sensing layer functions as an important integrity checkpoint, as sensors must process untrusted environmental signals in real-time. Optical spoofing (SP-A1) injects false LiDAR returns, while resonant ultrasound (SP-A3) destabilizes MEMS gyroscopes. Verifying the authenticity of external inputs in real time becomes difficult. Defenses like cross-modal verification (SP-D1) and coded illumination (SP-D2) leverage sensor redundancy, but the high volume and rate of sensor data make exhaustive validation impractical, leaving systems vulnerable to sophisticated physical manipulation techniques.
@@
Data Processing Layer — timing and fusion correctness. This layer is governed by strict real-time constraints (typically 1–10,ms). Even minor timing disruptions, such as from cache contention (DP-A1) or biased state estimation (DP-A4), can degrade system stability without triggering alarms. Defenses like timing guards (DP-D1) and memory-safe pipelines (DP-D2) mitigate such risks but often at a latency costs that real-time controllers cannot accommodate.
@@
Middleware Layer — Communication Infrastructure Attack Surface. As the backbone of communication, middleware enables high-throughput data exchange. Yet default configurations often favor usability over security. Topic spoofing (MW-A1) exploits missing authentication checks, while replay attacks (MW-A2) exploit missing freshness checks. Although cryptographic protections (MW-D1) and anomaly detection (MW-D2) can address these threats, their deployment is incomplete due to performance considerations, creating exploitable security gaps.
@@
Decision-Making Layer — Autonomous Control Under Adversarial Influence. As autonomy shifts toward ML-driven control, threats evolve from conventional faults to adversarial manipulation. Attacks exploit the opacity of AI models: adversarial inputs (DM-A1) mislead perception, while model poisoning (DM-A5) embeds latent backdoors. These attacks often avoid triggering system safeguards or error conditions, making detection particularly difficult. Defensive techniques like adversarial training (DM-D4) and certified inference (DM-D1) show promise but can be computationally costly. Compromises in this layer typically result in subtle misbehavior, where robots appear to function normally while executing attacker objectives.
@@
Application Layer — Remote Control Interface and Update Mechanism. As the interface for scripts, APIs, and OTA updates, this layer offers the most direct path to remote manipulation. Attacks are often semantic: script edits (AP-A1) silently alter behavior without memory corruption; unauthenticated APIs (AP-A2) enable full override. The main vulnerability lies in insufficient runtime policy enforcement. While defenses like signed code (AP-D1) and secure OTA mechanisms (AP-D3) exist, they are inconsistently deployed, allowing a breach here to bypass all lower layer protections.
@@
Social Interface Layer — Human-Robot Interaction as Attack Vector. This layer manages human-facing interactions through speech, vision, and gesture recognition systems. These modalities present unique attack surfaces, such as inaudible voice commands (SI-A1) or visual impersonation (SI-A3), which can trigger unauthorized actions without user awareness. Defending this inherently open interface is complex. Technical countermeasures like hardware filtering (SI-D1) and robust automated speech recognition (ASR) (SI-D2) help, but attacks often exploit social expectations rather than system flaws. Effective defenses must therefore combine technical safeguards with user trust boundary training (SI-D5) and privacy-first design principles.
@@ 5.1 The RISK-MAP Method
The RISK-MAP method is built on the set of 39 attack vectors $\mathcal{A}=\{a_{1},\ldots,a_{39}\}$ and 35 defense mechanisms $\mathcal{D}=\{d_{1},\ldots,d_{35}\}$ identified in our SoK. The methodology comprises four sequential steps: first establishing baseline threat severity across all identified attacks, then adjusting these assessments for platform-specific applicability, modeling the effectiveness of available defenses for the target platform, and finally aggregating these components into a comprehensive security score.
@@
First, we construct the baseline Coverage Matrix $\Gamma\in\mathbb{R}^{39\times 35}$. Each entry $\gamma_{ij}$ represents the inherent capability of defense $d_{j}$ to mitigate attack $a_{i}$, using the following discrete scale. These values are guided by empirical evidence of effectiveness in the literature.
@@
Second, we define a platform-specific Implementation Vector $\mu^{P}$, which captures the extent to which each defense $d_{J}$ is actually deployed on platform $P$. The implementation level is scored as:
@@
The Effective Coverage of defense $d_{j}$ against attack $a_{i}$ on platform $P$ is then the product of its baseline capability and its specific implementation:
$\varepsilon_{ij}^{P}=\gamma_{ij}\times\mu_{j}^{P}$
Step 4: Calculating Final Scores: To compute the combined effectiveness of multiple defenses against a single attack $a_{i}$, we assume that the defenses act independently. The total coverage for that attack on platform $P$ is:
$\kappa_{i}^{P}=1-\prod_{j=1}^{35}\left(1-\varepsilon_{ij}^{P}\right)$
@@
Layer-Specific Diagnostic Analysis: While the aggregated score is useful for high-level comparison, it can mask critical architectural weaknesses. To identify such localized weaknesses, we also compute a LayerScore for each of the seven architectural layers ($\ell$). This score measures how much a layer’s resident defenses contribute to mitigating all of the threats in the threat landscape. Let $\mathcal{D}_{\ell}\subset\mathcal{D}$ be the subset of defenses primarily associated with layer $\ell$. The combined coverage against attack $a_{i}$ using only these defenses is:
@@
These seven scores, scaled from 0-5, are presented in a chart (Figure 4(a)) to enable the visualization of asymmetric defense coverage and identification of layers where hardening would be beneficial.
Addressing Methodological Uncertainties with Monte Carlo Simulation: While our assessment process is structured and reproducible its inputs are subject to inherent uncertainty. Attack likelihoods, impact scores, and defense effectiveness ratings rely on a combination of historical data, evolving threat intelligence, system-specific details, and expert judgment, which can be incomplete or subjective. To address this limitation and ensure the robustness of our assessment results, we incorporate Monte Carlo simulation in our assessment [6, 64]. By repeatedly running the RISK-MAP assessment (1,000 iterations) with input values randomly sampled within a plausible range ($\pm 25\%$), we generate a distribution of possible security scores. This allows us to compute confidence intervals for our final results, providing a statistically grounded measure of how robust a platform’s security posture is despite the uncertainty of the underlying data.
@@ 5.2 Case Studies
To demonstrate RISK-MAP’s practical utility, we applied the framework on three representative humanoid platforms: Digit, G1 EDU, and Pepper. The assessment was based on publicly available datasheets, official documentation, and a standard deployment scenario. The results of the assessment are presented in To assist users in interpreting results, the assessment outputs are presented in multiple complementary formats of Figures 4(a)4(b) and Table 3.
@@
(b) Top-10 residual risks heatmap.
Figure 4: RISK-MAP security assessment results for three humanoid platforms.
Layer diagnostics. The radar chart (Figure 4(a)) provides a granular, layer-specific diagnostic for each platform, highlighting tactical and operational vulnerabilities across each of the three humanoids. This view enables engineering teams to prioritize targeted, subsystem-specific hardening efforts.
Residual-risk focus. The heatmap (Figure 4(b)) identifies the top-10 persistent attack vectors that remain after considering existing defenses. This offers a strategic view of cross-layer vulnerabilities that single-layer evaluations might miss. It assists security teams to address systemic risks in areas like perception integrity, memory safety, and AI model security.
@@ 6 Summary
Humanoids’ unique security challenges demand systematic analysis beyond conventional IT or CPS security models. This work addresses this gap through three contributions: a 7-layer security model organizing 39 attack vectors and 35 defense mechanisms across the humanoid stack, the RISK-MAP quantitative assessment method with Monte Carlo-based risk scoring enabling evidence-based security prioritization, and empirical evaluation of three deployed platforms revealing systematic vulnerabilities across all platforms and defense strategies prioritizing higher-layer protections while leaving fundamental attack vectors unaddressed. This work provides the first comprehensive, quantitatively grounded security model specifically designed for humanoids, establishing a foundation for proactive security assessment moving beyond reactive patching toward systematic threat mitigation. As these robots become increasingly autonomous and integrated into human environments, today’s security decisions will fundamentally determine their trustworthiness for decades—making this systematic approach essential for safe deployment.

### v2

@@ 4.2 A Layer-by-Layer Analysis of Threats and Defenses
Physical Layer - hardware, power, and safety baseline. The physical layer represents the most critical point of potential system compromise. Runtime attacks like firmware reflashing (P-A1) can embed persistent malware via debug interfaces, bypassing all software controls. Sensor bias (P-A2) distorts physical inputs, inducing silent drift in downstream systems. Since these attacks occur below the software stack, conventional detection mechanisms often miss them. This risk is amplified by perimeter-less operation; attackers need not breach an IT facility to reach the asset; they can target the robot in situ (casing, port access, sensor spoofing), then escalate across layers.
@@
Sensing and Perception Layer - Integrity Control Point for Environmental Input. The sensing layer functions as an important integrity checkpoint, as sensors must process untrusted environmental signals in real-time. Optical spoofing (SP-A1) injects false LiDAR returns, while resonant ultrasound (SP-A3) destabilizes MEMS gyroscopes. Verifying the authenticity of external inputs in real-time becomes difficult. Defenses like cross-modal verification (SP-D1) and coded illumination (SP-D2) leverage sensor redundancy, but the high volume and rate of sensor data make exhaustive validation impractical, leaving systems vulnerable to sophisticated physical manipulation techniques.
@@
Data Processing Layer - timing and fusion correctness. This layer is governed by strict real-time constraints (typically 1–10,ms). Even minor timing disruptions, such as from cache contention (DP-A1) or biased state estimation (DP-A4), can degrade system stability without triggering alarms. Defenses like timing guards (DP-D1) and memory-safe pipelines (DP-D2) mitigate such risks but often at a latency cost that real-time controllers cannot accommodate.
@@
Middleware Layer - Communication Infrastructure Attack Surface. As the backbone of communication, middleware enables high-throughput data exchange. Yet default configurations often favor usability over security. Topic spoofing (MW-A1) exploits missing authentication checks, while replay attacks (MW-A2) exploit missing freshness checks. Although cryptographic protections (MW-D1) and anomaly detection (MW-D2) can address these threats, their deployment is incomplete due to performance considerations, creating exploitable security gaps.
@@
Decision-Making Layer - Autonomous Control Under Adversarial Influence. As autonomy shifts toward ML-driven control, threats evolve from conventional faults to adversarial manipulation. Attacks exploit the opacity of AI models: adversarial inputs (DM-A1) mislead perception, while model poisoning (DM-A5) embeds latent backdoors. These attacks often avoid triggering system safeguards or error conditions, making detection particularly difficult. Defensive techniques like adversarial training (DM-D4) and certified inference (DM-D1) show promise but can be computationally costly. Compromises in this layer typically result in subtle misbehavior, where robots appear to function normally while executing attacker objectives.
@@
Application Layer - Remote Control Interface and Update Mechanism. As the interface for scripts, APIs, and OTA updates, this layer offers the most direct path to remote manipulation. Attacks are often semantic: script edits (AP-A1) silently alter behavior without memory corruption; unauthenticated APIs (AP-A2) enable full override. The main vulnerability lies in insufficient runtime policy enforcement. While defenses like signed code (AP-D1) and secure OTA mechanisms (AP-D3) exist, they are inconsistently deployed, allowing a breach here to bypass all lower layer protections.
@@
Social Interface Layer - Human-Robot Interaction as Attack Vector. This layer manages human-facing interactions through speech, vision, and gesture recognition systems. These modalities present unique attack surfaces, such as inaudible voice commands (SI-A1) or visual impersonation (SI-A3), which can trigger unauthorized actions without user awareness. Defending this inherently open interface is complex. Technical countermeasures like hardware filtering (SI-D1) and robust automated speech recognition (ASR) (SI-D2) help, but attacks often exploit social expectations rather than system flaws. Effective defenses must therefore combine technical safeguards with user trust boundary training (SI-D5) and privacy-first design principles.
@@ 5.1 The RISK-MAP Method
The RISK-MAP method is built on the set of 39 attack vectors $\mathcal{A}=\{a_{1},\ldots,a_{39}\}$ and 35 defense mechanisms $\mathcal{D}=\{d_{1},\ldots,d_{35}\}$ identified in our SoK. The method comprises four steps: first establishing baseline threat severity across all identified attacks, then adjusting these assessments for platform-specific applicability, modeling the effectiveness of available defenses for the target platform, and finally aggregating these components into a comprehensive security score.
@@
First, we construct the baseline Coverage Matrix $\Gamma\in\mathbb{R}^{39\times 35}$. Each entry $\gamma_{ij}$ represents the capability of defense $d_{j}$ to mitigate attack $a_{i}$, using the following scale. These values are guided by empirical evidence of effectiveness in the literature.
@@
Second,We define an Implementation Vector $\mu^{P}$ for each platform $P$ that measures how well each defense $d_{J}$ is deployed. The implementation level is scored as:
@@
The Effective Coverage ($\varepsilon_{ij}^{P}$) of defense $d_{j}$ against attack $a_{i}$ on platform $P$ is then the product of its baseline capability ( $\gamma_{ij}$) and its specific implementation ($\mu_{j}^{P}$):
$\text{Effective Coverage }\varepsilon_{ij}^{P}=\gamma_{ij}\times\mu_{j}^{P}$
Step 4: Calculating Final Scores: To compute the combined effectiveness of multiple defenses against a single attack $a_{i}$, we assume that the defenses act independently. The total coverage for that attack on platform $P$ ($\kappa_{i}^{P}$) is:
$\text{Total Coverage }\kappa_{i}^{P}=1-\prod_{j=1}^{35}\left(1-\varepsilon_{ij}^{P}\right)$
@@
Layer-Specific Diagnostic Analysis: While the aggregated score is useful for high-level comparison, it can mask critical architectural weaknesses. To identify such localized weaknesses, we also compute a LayerScore for each of the seven architectural layers ($\ell$). This score measures how much a layer’s resident defenses contribute to mitigating all of the threats in the threat landscape. Let $\mathcal{D}_{\ell}\subset\mathcal{D}$ be the subset of defenses primarily associated with layer $\ell$. The combined coverage ($\kappa_{i}^{P,\ell}$) against attack $a_{i}$ using only these defenses is:
@@
These 7 scores, scaled from 0-5, are presented in a chart (Figure 4(a)) to help in the visualization of asymmetric defenses and identification of layers where hardening would be beneficial.
Addressing Methodological Uncertainties with Monte Carlo Simulation: Our assessment inputs—attack likelihoods, impact scores, and defense effectiveness—involve uncertainty arising from historical data, evolving threat intelligence, system-specific details, and expert judgment. We addressed this by running 1,000 Monte Carlo [6, 69] iterations of RISK-MAP with input values randomly sampled within ($\pm 25\%$)intervals, providing statistically robust measures of platform security despite underlying data uncertainty.
@@ 5.2 Case Studies
To demonstrate RISK-MAP’s practical utility, we applied the framework on three representative humanoid platforms: Digit, G1 EDU, and Pepper. The assessment was based on publicly available datasheets, official documentation, and a standard deployment scenario. The results of the assessment are presented in Figures 4(a)4(b) and Table 3. To assist users in interpretation, we provide multiple complementary views.
@@
(b) Top-3 cross-layer vulnerability propagation
Figure 4: RISK-MAP results on three humanoids. (a) Layer coverage (seven layers; LayerScore 0–5). (b) Each row shows one robot’s highest-risk two-hop attack paths across the autonomy stack (P→SI). Line width encodes Cascade Residual Risk (CRR); line style indicates ranking. Recurring transitions reveal structural attack corridors; thickness variations expose robot-specific coverage gaps.
Layer diagnostics. The radar chart (Figure 4(a)) provides a layer-specific diagnostic for each platform, highlighting tactical and operational vulnerabilities across each of the three humanoids. This view enables engineering teams to prioritize targeted, subsystem-specific hardening efforts.
Cross-layer vulnerability propagation. We extend our RISK-MAP assessment to analyze cross-layer vulnerability propagation across our seven-layer model. Using coupling matrices derived from structural feasibility, empirical evidence, and mitigation strength, we enumerate two-hop attack cascades and compute Cascade Residual Risk (CRR) for each robot-path-attack combination. Figure 4(b) reveals systematic attack corridors, with DP→DM and S→DP transitions dominating high-risk paths across all robots, while robot-specific coverage gaps create varying CRR magnitudes for identical attack paths. Detailed methodology and mathematical formulation are provided in Appendix A.
@@ 6 Summary
This work addresses critical security challenges as humanoid robots transition to real-world deployment. Our systematic analysis of 39 attacks and 35 defenses across a seven-layer model demonstrates that current humanoid security is fundamentally inadequate for safe human-centered operation. Our RISK-MAP method reveals systemic vulnerabilities across 3 humanoids. While developers focus on application protections, critical foundational layers remain dangerously exposed. Cross-layer cascade analysis shows attackers can exploit these gaps for privilege escalation and safety bypasses. Monte Carlo validation across 1,000 trials confirms the robustness of our risk measurements. Humanoid systems exhibit higher vulnerability densities than mature cyber-physical systems, with multiple high-impact attacks cascading across multiple layers. As humanoids deploy in various industries, this security debt threatens public trust and safety. Our coupling matrix analysis reveals that defensive imbalances create systematic bottlenecks where single exploits can cascade into full system compromise. These architectural vulnerabilities require fundamental design changes rather than superficial patches. The robotics industry must transition from reactive patching to proactive defense architectures. Our methods enables practitioners to identify gaps, prioritize remediation, and validate improvements before deployment. The seven-layer model provides a standardized taxonomy for security assessment across diverse humanoid platforms. Immediate adoption of our methodology by manufacturers could prevent the security debt that has plagued other emerging technologies. As humanoids become ubiquitous in critical infrastructure, the window for establishing robust security foundations is rapidly closing and works such as ours can make the humanoid ecosystems to become safe and secure.
Appendix A Cross-Layer Vulnerability Propagation
To understand how security failures spread across humanoid robot systems, we build a $7\times 7$ matrix $D$ that maps how problems in one layer can affect other layers.
Matrix Construction: We model the coupling matrix as:
$D=(\alpha S+\beta E)\circ(1-M),$ (2)
  • $S\in[0,1]^{7\times 7}$: How easily problems could spread between layers based on system design,
  • $E\in[0,1]^{7\times 7}$: Real-world evidence of problems actually spreading this way,
  • $M\in[0,1]^{7\times 7}$: How well current security measures block this spread,
  • $\alpha,\beta\geq 0$ with $\alpha+\beta=1$: How much we trust design analysis ($\alpha$=0.6) vs. real incidents ($\beta$=0.4),
  • $\circ$: Combines values element by element.
We mark each layer as fully connected to itself $D_{ii}=1$ fbut don’t count these self-connections in our analysis.
Avoiding Double Counting
We separate two types of protection: edge mitigations $M$ that block spread between layers, and layer coverage $C_{r}(\ell)$ that protects within each layer. This prevents counting the same defense twice.
Finding Vulnerability Paths
We look for two-step attack paths: $\pi=\ell_{i}\rightarrow\ell_{j}\rightarrow\ell_{k}$ where each hop meets a threshold and the full chain exceeds minimum propagation strength.
Per-hop thresholds:
$D_{ij}\geq\varepsilon,\qquad D_{jk}\geq\varepsilon.$ (3)
Path strength:
$P(\pi)=D_{ij}\cdot D_{jk}\;\;\geq\;\;\text{MIN\_PROP}.$ (4)
Parameters Unless otherwise stated, we use:
$\alpha=0.6,\quad\beta=0.4,\quad\varepsilon=0.3,\quad\text{MIN\_PROP}=0.1.$
We tested longer 3-step paths but found that they didn’t change which paths were most dangerous. Risk Metrics
Cascade Risk (CRR) For robot $r$ and attack $a$, the Cascade Risk along chain $\pi$ indicates residual risk after going through a path and encountering defenses.
$\operatorname{CRR}(\pi,r,a)=\operatorname{clip}\!\left(\,P(\pi)\cdot w_{a}\cdot U_{r}(\pi)\,\right),$ (5)
  • $w_{a}$: Attack weight (based on impact, likelihood, detectability),
  • $U_{r}(\pi)=\prod_{\ell\in\pi}\bigl{(}1-C_{r}(\ell)\bigr{)}$: Multiplicative defense gap for robot $r$ along $\pi$,
  • $C_{r}(\ell)\in[0,1]$: Per-layer coverage of defenses at layer $\ell$ for robot $r$,
  • $\operatorname{clip}\!\left(x\right)=\min(1,\max(0,x))$: Clipping function.
Cascade Containment Index (CCI) The complementary metric indicates how well the system contains the attack.
$\operatorname{CCI}(\pi,r,a)=1-\operatorname{CRR}(\pi,r,a).$ (6)
Interpretation.
  • High CRR: Strong propagation paths with high-weight attacks meeting weak defenses.
  • High CCI: Effective containment through edge mitigations and layer coverage.
Uncertainty Analysis
We assess robustness via Monte Carlo simulation with 1,0001{,}000 runs, applying i.i.d. $\pm 25\%$ uniform noise to:
  • $\gamma$ (defense efficacy),
  • $\mu$ (deployment level),
  • $w_{a}$ (attack weights).
Results report median values with 5th/95th percentile confidence bands.
Key Findings
Structural Attack Corridors The cascades highlight recurring structural paths, especially $P\!\rightarrow\!S\!\rightarrow\!DP$, $S\!\rightarrow\!DP\!\rightarrow\!DM$ and $DP\!\rightarrow\!DM\!\rightarrow\!AP$ . These corridors consistently rank among the most critical across robots. Their recurrence indicates that vulnerabilities concentrate at these layer junctions. Strengthening these choke-points would yield disproportionate security improvements.
Cross-Robot Risk Heterogeneity While the same structural paths appear across Digit, G1_EDU, and Pepper, their Chain Risk Ratings (CRR) differ noticeably. For example, Digit shows higher CRRs on DP $\rightarrow$ DM $\rightarrow$ AP (CRR $\approx 0.113$), whereas Pepper exhibits comparable risk in S $\rightarrow$ DP $\rightarrow$ DM (CRR $\approx 0.078$). This heterogeneity stems from differences in layer coverage and implementation choices, highlighting that robot-specific defenses are necessary even when structural paths overlap.
Attack Path Dominance The results confirm that short, high-leverage paths (h=2) dominate the Top-3 cascades, often outranking longer chains. For instance, Digit’s DP $\rightarrow$ DM $\rightarrow$ AP (CRR =0.113=0.113) is more severe than longer cascades not captured in the Top-3.